OpenStudy (aaronandyson):

If (3x+1):(5x+3) is the triplicate ratio of 3:4,find the value of x

OpenStudy (aaronandyson):

@chmvijay

OpenStudy (cathyangs):

This can be rewritten as: \[\frac{ (3x+1) }{ (5x+3)} = \frac{ 3 }{ 4 }\] Can you use cross-products to get this from fraction form to fraction-less equations? Then you can perform the same mathematical functions ( +, -, *, /) to each side to keep the two sides equal. Solve to get the form x = ? .

OpenStudy (aaronandyson):

shouldn't 3 and 4 be 64 ans 27 as their respective cubes..??

OpenStudy (cathyangs):

Actually, I am corrected. I just Google-ed "triplicate ratio" and I think this means that the cubes of each binomial are in the ratio 3 to 4. ( http://www.thefreedictionary.com/Triplicate+ratio)

OpenStudy (aaronandyson):

triplicate ratio of a:b is a^3 : b^3

OpenStudy (cathyangs):

Or...the wording suggests that it's the triplicate ratio of 3 to 4? (27 to 64 as you said). Sorry about that! Then the proportion would be: \[\frac{ 3x+1 }{ 5x + 3} = (\frac{ 3 }{ 4 })^{3} = \frac{ 27 }{ 64 }\]

OpenStudy (cathyangs):

What do you get as your answer for x ? I can tell if it matches what I got.

OpenStudy (aaronandyson):

9/11

OpenStudy (cathyangs):

I don't get that as my answer. As an intermediate, I got: 192 x + 64 = 135 x + 81, from cross-multiplying. Then I used algebra to simplify to get x = ? (also verified my answer, so it's correct). What do you get as your new answer?

OpenStudy (aaronandyson):

57x = 17 x = 17/57

OpenStudy (cathyangs):

That's correct! Good work! :) Don't forget to close the question.

imqwerty (imqwerty):

yes x=17/57 is correct